OpenStudy (anonymous):

if sin^-1(1/3) + cos^-1(x) = pi/2, then find x.

OpenStudy (aravindg):

OK now I can help :)

OpenStudy (aravindg):

first of all tell me is 90 in degrees?

OpenStudy (aravindg):

oh you edited it ok

OpenStudy (aravindg):

you see we have a fundamental property that \[\sin^{-1} x+ \cos^{-1} x =\dfrac{\pi}{2}\] Use this idea to get your answer

OpenStudy (anonymous):

ok thank u

OpenStudy (aravindg):

yw :)